Calculate the approximate value of √26 to four decimal places using Taylor series method.


1
Expert's answer
2021-07-26T16:55:49-0400

Taylor series is given as

"f(a+h)=f(a)+h.f'(a)+{h^2\over 2!}f''(a)+{h^3\over 3!}f'''(a)+..."

Let's Rewrite "26" in the following form:

"26=25+1" Since we know square root of "25"

Now, let's have a function that will enable us use Taylor series method

"f(26)=f(25+1)"

"\implies" "a=25" and "h=1"

Since we are asked to find the square root

"\implies f(a)=a^{1\over 2} \implies f(a+1)=\sqrt {26}"

Substituting the value of "a" we have

"f(a)=25^{1\over 2} =5"

"f'(a)={1\over 2}a^{{1\over 2}-1}={1\over 2}a^{-{1\over 2}}={{1\over 2}\over a^{1\over2}}={1\over 2a^{1\over 2}}"

Substituting the value of "a" we have

"f'(a)={1\over 2(25)^{1\over 2}}=0.1"

"f''(a)=(-{1\over 2})({1\over 2})a^{-{1\over 2}-1}=-{1\over 4}a^{-{3\over 2}}={-{1\over 4}\over a^{3\over 2}}=-{1\over 4a^{3\over 2}}"

Substituting the value of "a" we have

"f''(a)=-{1\over 4(25)^{3\over 2}}=-0.002"

"f'''(a)=({-{3\over 2}})({-{1\over 4}})a^{-{3\over 2}-1}={3\over 8}a^{-{5\over 2}}={{3\over 8}\over a^{5\over 2}}={3\over 8a^{5\over 2}}"

Substituting the value of "a" we have

"f'''(a)={3\over 8(25)^{5\over 2}}=0.00012"

From Taylor series method, we have

"\sqrt {26}=f(a+1)=f(a)+1.f'(a)+{1^2\over 2!}f''(a)+{1^3\over 3!}f'''(a)+..."

"\implies \sqrt {26}= f(a+1)=f(a)+f'(a)+{f''(a)\over 2}+{f'''(a)\over 3!}+..."

But

"f(a)=5" , "f'(a)=0.1" , "f''(a)=-0.002" , "f'''(a)=0.00012"

"\implies \sqrt {26}= f(a+1)=5+0.1+{(-0.002)\over 2}+{0.00012\over 3!}=5.09902"

"\therefore" "\sqrt{26}=5.0990" correct to four decimal places